# Detecting Multimedia Generated by Large AI Models: A Survey

### v1

@@ 3.5.1 Pure Detection
$\clubsuit$ Text-assisted. This type of method utilizes text-image or prompt-image pair as input and focuses on extracting and learning features from both text and images to improve DM-generated image detection, see Fig. 10 a). Amoroso et al. [151] leverage the semantic content of textual descriptions alongside visual data. They introduce a contrastive-based disentangling strategy to analyze the role of the semantics of textual descriptions and low-level perceptual cues. The experiments are conducted on their proposed COCOFake dataset, where CLIP [138] is employed as the backbone for feature extraction. These extracted features are then used to train a logistic regression model. They find that the visual features extracted from the generated image still retain the semantic information of the original caption used to create it, which allows them to distinguish between natural images and the generated ones using only semantic cues while neglecting the perceptual ones.
$\clubsuit$ Text-image Inconsistency. This refers to a type of misinformation where the textual content does not align with or accurately represent the original meaning or intent of the associated image. This type of misinformation is commonly presented in the guise of news [230] to mislead the public, see Fig. 10 b). Tan et al. [13] create a NeuralNews dataset which contains articles with LLM-manipulated text and authentic images tailored for such task. They further propose DIDAN (Detecting Cross-Modal Inconsistency to Defend Against Neural Fake News) framework that effectively leverages visual-semantic inconsistencies between article text, images, and captions, offering a novel approach to counteract neural fake news.
@@ 3.5.2 Beyond Detection
$\clubsuit$ Attribution. This refers to research works that utilize a multimodal learning approach, combining prompt information with image features, to enhance the accuracy of attributing DM-generated images to their source model, shown in Fig. 11 a). Sha et al. [150] build an image-only detector (ResNet18) and a hybrid detector (take advantage of CLIP’s [138] image and text encoders for feature extraction). Empirical results show that learning images with prompts achieves better performance than image-only attribution regardless of the dataset.
@@
$\clubsuit$ Interpretability. It provides evidence with detection results by leveraging multimodal media, shown in Fig. 11 c). To combat misinformation and emphasize the significance of scalability and explainability in misinformation detection, Xu et al. [147] propose a conceptual multi-modal framework composed of 4 levels: signal, perceptual, semantic, and human, for explainable multimodal misinformation detection.
@@
$\clubsuit$ Empirical Study It provides insights into the feasibility of detecting generated images by leveraging multimodal media shown in Fig. 11 e). [143] shows that it is possible to detect the generated images using simple Multi-Layer Perceptrons (MLPs), starting from features extracted by CLIP [138] or traditional CNNs. They found that incorporating the associated textual information with the images rarely leads to significant improvement in detection results but that the type of subject depicted in the image can significantly impact performance. Papadopoulos et al. [144] highlight that specific patterns and biases in multimodal misinformation detection benchmarks can result in biased or unimodal models outperforming their multimodal counterparts on a multimodal task. Therefore they create a ”VERification of Image-TExt pairs” (VERITE) benchmark to address the unimodal bias in multimodal misinformation detection.
4 Tools
@@
In this section, we present a selection of popular, user-friendly detection tools we have evaluated, see Table II.
@@
$\clubsuit$ Text. In the domain of LLM-generated text detection tools, diverse instances illuminate specific strengths and drawbacks. Copyleaks [232] offers high accuracy across 30 languages but faces performance delays and lacks text exclusion options. ZeroGPT [233] is user-friendly but doesn’t support file analysis. Though Winston AI[234] presents a free OCR (Optical Character Recognition)-powered alternative with an accessible interface, it needs enhancements in reporting and language support. Simplicity and speed take center stage with Crossplag’s[235], but it is confined to English and lacks graphical elements. GLTR[236] delivers in-depth, color-coded text analysis but is less intuitive for beginners and struggles with newer technologies. Content at Scale[237] yields easy-to-interpret results but may exhibit inconsistencies in AI rewrites. Undetectable.ai [238] supports various models but may discourage users due to pricing concerns. Illuminarty[239] excels in detecting AI-generated texts but introduces complexity and slower processing for basic needs. Finally, Is it AI[240] matches other tools in function but faces reliability issues, especially with short or non-English texts.
@@
Modality Tool Company Link Reference
@@
AI Content Detector Copyleaks Link [232]
@@
AI Content Detector, ChatGPT detector ZeroGPT Link [233]
@@
AI Content Detector Winston AI Link [234]
@@
AI Content Detector Crossplag Link [235]
@@
Giant Language model Test Room GLTR Link [236]
The AI Detector Content at Scale Link [237]
AI Checker Originality ai Link [241]
@@
Advanced AI Detector and Humanizer Undetectable ai Link [238]
@@
AI Content Detector Writer Link [242]
@@
AI Content Detector Conch Link [243]
Illuminarty Text Illuminarty Link [239]
@@
Text AI-Generated Text Detector Is it AI Link [240]
AI or Not image AI or Not Link [244]
@@
AI-Generated Image Detector Is it AI Link [245]
Illuminarty Image Illuminarty Link [246]
SynthID Google Link [247]
Advanced AI Image Detector Content at Scale Link [248]
Image AI Image Detector Huggingface Link [249]
Audio AI or Not audio AI or Not Link [244]
TABLE II: Existing popular detection tools for exposing LAIM-generated multimedia.
@@
$\clubsuit$ Image. In evaluating tools for detecting images generated by DMs, AI or Not [244] demonstrated fast execution while identifying fake images and supported diverse formats. However, it cannot provide stable results. In addition, Is it AI[245] and Content at Scale[248] offer a user-friendly interface with limitations in detecting certain content types. Illuminarty[246] provides in-depth analysis but faces usability complexities, and Huggingface[249] excelled in accessibility with limitations in advanced analysis. Google has developed SynthID[247], specifically designed for images generated by Imagen[250].
$\clubsuit$ Audio. AI or Not[244] also provides an AI audio detection tool which sets it apart from other AI detection tool product lineups. Currently, they support WAV, MP3, and FLAC audio files of a maximum size of 100MB and have a duration of more than 5 seconds.

### v4

@@ 3.4.1 Pure Detection
3.4.2 Beyond Detection
$\clubsuit$ Generalization. Ren et al.[240] propose a disentanglement-based framework that targets domain-agnostic vocoder artifacts for generalization. Their model leverages a dual encoder architecture, separating content and artifact features, paired with multi-task and contrastive learning to classify domain-specific and domain-agnostic features. Their method improves detection across both seen and unseen vocoders.
3.4.3 Analysis
$\clubsuit$ Insights. Though research on detecting LAIM-generated audio is still emerging, it reveals key patterns distinguishing synthetic from human speech. Neural vocoders often leave detectable artifacts, such as unnatural harmonics or frequency discontinuities, in spectrotemporal features [132]. However, detectors trained on specific vocoders often overfit, limiting generalization to unseen methods. To address this, recent studies propose disentangling content from vocoder-specific artifacts to capture domain-agnostic cues [240]. Incorporating vocoder classification into multitask learning further encourages models to prioritize forensic over linguistic or speaker features.
$\clubsuit$ Recommendations. Disentangling content from vocoder-specific artifacts, using dual-encoder architectures and contrastive learning, can improve generalization across unseen synthesis models [240]. Multitask learning that jointly performs vocoder classification and authenticity detection has shown promise, as it encourages feature extractors to focus on generation specific anomalies while remaining robust to linguistic variation [132]. Additionally, leveraging spectrogram-based representations enriched with perceptually motivated features (e.g., phase, pitch, or formant structure) can help detect subtle inconsistencies introduced by synthesis pipelines. Robust evaluation should further consider compression artifacts and language diversity, as both factors significantly affect detection reliability. Building comprehensive, multilingual, and vocoder-diverse benchmarks will be essential to support the development and evaluation of generalizable audio detection systems.
$\clubsuit$ Social Media. A benchmark conducted by Müller et al. [131] shows that detectors that perform well on curated datasets like ASVspoof fail dramatically when faced with real-world, social media processed audio, experiencing up to a tenfold increase in error rates. Social media introduces platform specific degradations such as compression, resampling, and added noise, which obscure the synthetic artifacts that most current detectors rely on. Models using fixed-length inputs and limited feature representations, such as melspectrograms, were particularly fragile, while raw waveform-based models showed relatively better, though still insufficient, resilience.
@@ 3.5.1 Pure Detection
$\clubsuit$ Prompt-guided. This category includes methods that leverage prompts, instructions, or text-image pairs to guide the learning process of image or video forgery detection models. These approaches typically extract both visual and semantic features, using the language modality to improve performance in DM-generated media detection, as shown in Fig. 10 a). Amoroso et al. [145] leverage the semantic content of textual descriptions alongside visual data. They introduce a contrastive-based disentangling strategy to analyze the role of the semantics of textual descriptions and low-level perceptual cues. The experiments are conducted on their proposed COCOFake dataset, where CLIP [212] is employed as the backbone for feature extraction. These extracted features are then used to train a logistic regression model. They find that the visual features extracted from the generated image still retain the semantic information of the original caption used to create it, which allows them to distinguish between natural images and the generated ones using only semantic cues while neglecting the perceptual ones. Song et al. [120] propose MM-Det, a diffusion-generated video detector that learns a Multi-Modal Forgery Representation from large vision-language models via instruction tuning. Alongside the MMFR, the authors introduce an In-and-Across Frame Attention mechanism that captures spatial artifacts and temporal inconsistencies across frames. A dynamic fusion strategy adaptively weights both streams for final prediction. Bohacek et al. [239] propose Human Action CLIPS, a detection method targeting AI-generated human motion videos. Instead of relying on low-level artifacts, their approach uses multi-modal semantic embeddings (e.g., CLIP, SigLIP) combined with lightweight classifiers. Evaluated on their proposed DeepAction dataset, which includes videos from seven AI models and real footage, their method achieves robust detection even after compression.
$\clubsuit$ Text-image Inconsistency. This refers to a type of misinformation where the textual content does not align with or accurately represent the original meaning or intent of the associated image. This type of misinformation is commonly presented in the guise of news [222] to mislead the public, see Fig. 10 b). Tan et al. [13] create a NeuralNews dataset which contains articles with LLM-manipulated text and authentic images tailored for such task. They further propose DIDAN (Detecting Cross-Modal Inconsistency to Defend Against Neural Fake News) framework that effectively leverages visual-semantic inconsistencies between article text, images, and captions, offering a novel approach to counteract neural fake news. Huang et al. [238] introduces D-TIIL (Diffusion-based Text-Image Inconsistency Localization), which employs text-to-image diffusion models to localize semantic inconsistencies in text and image pairs.
@@ 3.5.2 Beyond Detection
$\clubsuit$ Attribution. This refers to research works that utilize a multimodal learning approach, combining prompt information with image features, to enhance the accuracy of attributing DM-generated images to their source model, shown in Fig. 11 a). Sha et al. [236] build an image-only detector (ResNet18) and a hybrid detector (take advantage of CLIP’s [212] image and text encoders for feature extraction). Empirical results show that learning images with prompts achieves better performance than image-only attribution regardless of the dataset. Building upon this, Keita et al. [237] introduced FIDAVL, a multitask framework leveraging vision-language models and soft prompt-tuning to jointly detect and attribute synthetic images to their originating generative models.
@@
$\clubsuit$ Interpretability. It provides evidence with detection results by leveraging multimodal media, shown in Fig. 11 c). To combat misinformation and emphasize the significance of scalability and explainability in misinformation detection, Xu et al. [231] propose a conceptual multi-modal framework composed of 4 levels: signal, perceptual, semantic, and human, for explainable multimodal misinformation detection. Huang et al. [232] introduce FFAA, a face forgery analysis assistant that redefines detection as an Open-World Face Forgery VQA task. It fine-tunes MLLMs (Multimodal Large Language Models) using hypothetical prompts and introduces a Multi-answer Intelligent Decision System to disambiguate difficult cases by reasoning over multiple candidate answers. FFAA not only provides fine-grained reasoning (e.g., forgery type, difficulty) but also outperforms prior detectors across six benchmark datasets. Chen et al. [233] propose X²-DFD, a framework that identifies strong and weak forgery-related features—such as facial structure and blending artifacts—through a feature assessment module. It enhances detection by combining prompt-based fine-tuning with external detectors, and goes beyond image-level predictions by providing feature-grounded textual explanations, thereby improving the transparency of detection results.
@@
MLLM-based. While existing multimodal localization methods often rely on spatial or frequency alignment across image-text pairs, recent advances leverage MLLMs (Multimodal Large Language Models) to enable localization with accompanying textual explanations. Xu et al.[227] introduce FakeShield, a pioneering framework that combines MLLMs with pixel-level segmentation and language-based justification. It produces both textual descriptions and tampering masks using a Domain Tag-guided Explainable Forgery Detection Module and a SAM-based localization module. Liu et al.[228] present ForgeryGPT, which supports multi-turn dialogue for forgery analysis by integrating a Mask-Aware Forgery Extractor into a customized LLM. It employs an Object-agnostic Forgery Prompt and a Vocabulary-enhanced Vision Encoder to capture multi-scale forgery patterns and semantic-level inconsistencies, enabling precise localization and rich textual explanation.
@@
$\clubsuit$ Empirical Study. It provides insights into the feasibility of detecting generated images by leveraging multimodal media shown in Fig. 11 e). [223] shows that it is possible to detect the generated images using simple Multi-Layer Perceptrons (MLPs), starting from features extracted by CLIP [212] or traditional CNNs. They found that incorporating the associated textual information with the images rarely leads to significant improvement in detection results but that the type of subject depicted in the image can significantly impact performance. Similarly,  [225] explores the effectiveness of CLIP with four distinct transfer learning strategies, including fine-tuning, linear probing, prompt tuning, and training an adapter network. They reveal that prompt tuning strategy which leverages both the image and text components of CLIP consistently achieves better performance compared with baselines and SOTA methods (e.g., Corvi et al. [102], Ojha et al. [98] Papadopoulos et al. [224] highlight that specific patterns and biases in multimodal misinformation detection benchmarks can result in biased or unimodal models outperforming their multimodal counterparts on a multimodal task. Therefore they create a ”VERification of Image-TExt pairs” (VERITE) benchmark to address the unimodal bias in multimodal misinformation detection. Jia et al. [226] investigate the capabilities of multimodal large language models (LLMs) in DeepFake detection. They input fake and real images with different prompts that embody the instruction and request to the LLM (e.g., GPT4) to detect DeepFake faces. They find that LLMs can achieve a notable AUC score of 75%. Despite their promise, these models face challenges in accurately recognizing authentic images and require sophisticated prompting techniques for optimal performance, all without depending on signal cues or traditional data-driven detection methods.
3.5.3 Analysis
$\clubsuit$ Insights. Multimodal detection methods capitalize on the interplay between visual, linguistic, and occasionally auditory signals to identify inconsistencies or artifacts that may not be evident within a single modality. A key insight is that semantic misalignment—such as mismatched captions or fabricated prompts—serves as a strong cross-modal signal for detection [13, 145]. Prompt-guided methods, contrastive frameworks, and visual question answering (VQA)-based systems leverage this to enhance both detection and attribution [235, 120, 236]. However, empirical studies caution that naive use of language features often yields marginal gains unless modality interaction is explicitly modeled [223, 224]. Moreover, advances in MLLMs introduce new opportunities for interpretability and localization, enabling models to generate rationales and segment manipulated regions simultaneously [232, 227].
These developments underscore the growing feasibility of multimodal detection systems in practical settings. Recent works such as FakeShield [227], ForgeryGPT [228], and FFAA [232] demonstrate that MLLMs can be effectively fine-tuned or guided to perform both detection and explanation tasks. These systems support multi-turn reasoning, open-ended prompts, and produce interpretable outputs like textual justifications and pixel-level mask, making them adaptable for detecting complex, real-world manipulations. As such, the feasibility of deploying MLLM-based detectors is supported by strong empirical evidence, especially in applications where scalable, explainable, and cross-modal understanding is essential. Despite this progress, challenges remain: multimodal detectors are susceptible to unimodal biases, often over-relying on dominant modalities, and performance varies significantly with image subject type or text domain [224, 211].
$\clubsuit$ Recommendations. To improve the robustness and effectiveness of multimodal detection, several directions are promising. First, enhancing cross-modal alignment through fine-grained contrastive learning can better capture subtle inconsistencies across modalities. To mitigate unimodal bias, training with balanced and semantically diverse data, alongside attention regularization, ensures the model does not over-rely on a dominant modality. Soft prompt tuning and instruction-guided learning can further support generalization to unseen domains by leveraging linguistic context as a guiding signal. Integrating detection and localization within unified transformer-based frameworks improves interpretability and supports pixel-level forensics. Additionally, incorporating MLLMs enables explainable detection through natural language rationales, though care must be taken to avoid hallucinated explanations. Finally, hybrid architectures combining spatial, frequency, and semantic features, along with well-constructed benchmarks such as DGM4 [14] and MMFakeBench [155], are critical to evaluating detectors under realistic, cross-modal manipulation scenarios.
$\clubsuit$ Social Media. In addressing the challenge of detecting social-media-processed AI-generated multimodal content, MMFakeBench [155] introduces MMD-Agent, a specialized detection model designed to tackle mixed-source misinformation. MMD-Agent decomposes the detection task into three complementary branches—textual, visual, and cross-modal inconsistency reasoning—leveraging large vision-language models (LVLMs) to handle the noisy, stylized, and semantically complex content commonly found on social platforms. By explicitly modeling reasoning across modalities and training on mixed-source forgeries, MMD-Agent achieves more robust performance under real-world distortions compared to traditional artifact-based models. The ILLUSION benchmark [158] reveals significant performance degradation when detectors are exposed to social-media-sourced deepfakes through extensive benchmarking of state-of-the-art unimodal and multimodal models, highlighting critical gaps in cross-modal generalization, robustness to noise, and zero-shot generalizability.
4 Tools and Evaluation Metrics
This section reviews practical tools and standard evaluation metrics commonly used in detecting LAIM-generated content.
4.1 Tools
@@
In this section, we present a selection of popular, user-friendly detection tools we have evaluated, see Table II. Several tools discussed in the multimodal section offer support for image, video, and audio detection. In the modality-specific subsections (image, video, and audio), we focus on tools that are specifically designed for a single modality.
@@
$\clubsuit$ Text. In the domain of LLM-generated text detection tools, diverse instances illuminate specific strengths and drawbacks. Copyleaks [360] offers high accuracy across 30 languages but faces performance delays and lacks text exclusion options. ZeroGPT [361] is user-friendly but doesn’t support file analysis. Though Winston AI[362] presents a free OCR (Optical Character Recognition)-powered alternative with an accessible interface, it needs enhancements in reporting and language support. Simplicity and speed take center stage with Crossplag’s[363], but it is confined to English and lacks graphical elements. GLTR[364] delivers in-depth, color-coded text analysis but is less intuitive for beginners and struggles with newer technologies. Brandwell [365] yields easy-to-interpret results but may exhibit inconsistencies in AI rewrites. Undetectable.ai [366] supports various models but may discourage users due to pricing concerns. Illuminarty[367] excels in detecting AI-generated texts but introduces complexity and slower processing for basic needs. Finally, Is it AI[368] matches other tools in function but faces reliability issues, especially with short or non-English texts.
TABLE II: Summary of existing popular detection tools for exposing LAIM-generated multimedia. The ‘Multi-platform’ label in the Type column indicates that these tools are not limited to web applications but can also be used or deployed on other platforms, such as mobile apps or Chrome extensions.
@@
Modality Tool Company Link Reference Type Open Source Cost
@@
AI Content Detector Copyleaks Link [360] Webapp & API Limited free usage
@@
AI Content Detector, ChatGPT detector ZeroGPT Link [361] Webapp & API Free usage
AI Detector GPTZero Link [369] Multi-platform Limited free usage
@@
AI Content Detector Winston AI Link [362] Webapp & API Limited free usage
@@
AI Content Detector Crossplag Link [363] Webapp Limited free usage
@@
Giant Language model Test Room GLTR Link [364] Webapp Free usage
The AI Detector Brandwell Link [365] Webapp Limited free usage
AI Checker Originality ai Link [370] Webapp & API Limited free usage
@@
Advanced AI Detector and Humanizer Undetectable ai Link [366] Webapp & API Limited free usage
@@
AI Content Detector Writer Link [371] Webapp & API Limited free usage
@@
AI Content Detector Conch Link [372] Webapp Limited free usage
Illuminarty Text Illuminarty Link [367] Webapp & API Limited free usage
@@
Text AI-Generated Text Detector Is it AI Link [368] Webapp & API Limited free usage
Liveness Detection, Facial Recognition Incode Link [373] Multi-platform Paid
AI or Not image AI or Not Link [374] Webapp & API Limited free usage
@@
AI-Generated Image Detector Is it AI Link [375] Webapp & API Limited free usage
Illuminarty Image Illuminarty Link [376] Webapp & API Limited free usage
AI Image Detector Undetectable ai Link [366] Webapp & API Limited free usage
SynthID Google Link [377] Webapp Limited free usage
The AI image detector Winston Link [362] Webapp & API Limited free usage
Image Advanced AI Image Detector Brandwell Link [378] Webapp Limited free usage
Deepware Scanner Deepware Link [379] Webapp & API Free usage
Video Attestiv Deepfake Video Detection Attestiv Link [380] Webapp & API Limited free usage
Pulse Inspect Pindrop Link [381] Multi-platform Paid
AI Voice Detector AI Voice Detector Link [382] Webapp & API Limited free usage
AI Speech Classifier ElevenLabs Link [383] Webapp & API Limited free usage
Audio AI or Not audio AI or Not Link [374] Multi-platform Limited free usage
Video, Image, and Audio Detector Deep Media Link [384] Multi-platform Limited free usage
Deepfake Detection Sensity AI Link [385] Multi-platform Paid
Hive AI’s Deepfake Detection API Hive AI Link [386] API Limited free usage
Resemble Detect Resemble AI Link [387] Webapp & API Limited free usage
DuckDuckGoose AI (Phocus) DuckDuckGoose AI Link [388] Webapp Paid
Sentinel Sentinel Link [389] Webapp & API Paid
Deepfake Detector Deepfake Detector Link [390] Multi-platform Limited free usage
DeepFake-o-meter U of Buffalo Link [391] Webapp Free usage
BioID BioID Link [392] Webapp & API Limited free usage
Get Real Protect Get Real Link [393] Multi-platform Paid
Multi-modal Reality Defender Reality Defender Link [394] Multi-platform Paid
TABLE III: Summary of evaluation metrics used for detecting LAIM-generated multimedia.
Modality Task Metric Description
Common Binary Classification AUC [43] Area under the ROC curve, measures discriminative ability.
F1-score [41] Harmonic mean of precision and recall.
FPR [48] Proportion of real instance incorrectly classified as generated.
Accuracy [60] Proportion of correctly classified instances.
Precision [14] Proportion of correctly identified generated instances among all identified generated instances.
Recall [166] Proportion of correctly identified generated instances among all actual generated instances.
Average Precision [254] The area under the precision-recall curve, summarizing precision over all recall levels as the threshold varies
Text Binary Classification AvgRec [34] Average of recall on human-written and LLM-generated texts.
Image Localization IoU (mIoU) [254] Intersection over Union (mean IoU).
Pixel-level AUC [92] Area Under the ROC Curve computed at the pixel level.
Pixel-level F1 [92] Harmonic mean of pixel-wise precision and recall.
PBCA [254] Pixel-wise Binary Classification Accuracy
Fairness Demographic Parity $F_{DP}$ [116] Measures the maximum difference in prediction rates across all demographic groups.
Max Equalized Odds $F_{MEO}$ [116] Captures the largest disparity in prediction outcomes (TPR or FPR) between any two demographic groups.
Equal Odds $F_{EO}$ [116] Measures the disparity in TPR and FPR between each subgroup and the overall population.
Overall Accuracy Equality $F_{OAE}$ [116] Measures the maximum accuracy gap across all demographic groups.
Individual Fairness $F_{IND}$ [116] Measures if similar individuals receive similar model outputs.
Video Binary Classification Video-level AUC [120] Area Under the ROC Curve computed at the video level.
Audio Binary Classification EER [132] Equal Error Rate. The point on the ROC curve where the false positive rate (FPR) equals the false negative rate (FNR).
t-DCF [240] Tandem Detection Cost Function. Evaluate the trade-off between detection errors and the cost of those errors.
Multimodal Image Grounding mIoU [92] Measures average overlap between predicted and true bounding boxes.
Text Grounding Token-level Precision [14] Percentage of correctly predicted manipulated tokens among those predicted as manipulated.
Token-level Recall [14] Percentage of correctly predicted manipulated tokens among all ground truth manipulated tokens.
Token-level F1 [14] Harmonic mean of precision and recall for manipulated token detection.
Text Explanation CSS [227]
Cosine Semantic Similarity. Assess the similarity between the predicted text and ground truth text
by calculating the cosine similarity between their high-dimensional semantic vectors.
ROUGE [228] Recall-Oriented Understudy for Gisting Evaluation. Assess the overall accuracy of the generated text explanations.
@@
$\clubsuit$ Image. In evaluating tools for detecting images generated by DMs, AI or Not [374] demonstrated fast execution while identifying fake images and supported diverse formats. However, it cannot provide stable results. In addition, Is it AI[375] and Brandwell [378] offer a user-friendly interface with limitations in detecting certain content types. Illuminarty[376] provides in-depth analysis but faces usability complexities, and Huggingface[395] excelled in accessibility with limitations in advanced analysis. Google has developed SynthID[377], specifically designed for images generated by Imagen[396].
$\clubsuit$ Video. Deepware Scanner [379] is a web-based platform designed for quick and accessible deepfake analysis. It allows users to link videos, particularly from platforms like YouTube, Twitter, and Facebook, and provides a probability score indicating the likelihood of manipulation. Aimed at general users such as journalists or content moderators, it prioritizes ease of use over technical depth. Attestiv [380] targets enterprise use cases with an AI-driven framework offering forensic media analysis, manipulation localization, and verification via digital fingerprinting. It provides a detailed Suspicion Score and supports API integration, making it suitable for industries like insurance, finance, and cybersecurity. While Deepware emphasizes usability, Attestiv offers greater analytical depth and system-level scalability.
$\clubsuit$ Audio. Commercial tools for detecting AI-generated audio are newer and often packaged as enterprise fraud solutions. Pindrop’s Pulse Inspect is one prominent example: it is a web-based tool for detecting AI-generated speech in any digital audio or video file with what it claims is a significantly high degree of accuracy: 99% [397]. AI Voice Detector [382] lets users upload audio files and analyzes them to ascertain the likelihood of AI generation. Key features include background noise reduction to enhance detection accuracy and a browser extension for real-time analysis of audio content on platforms like YouTube and Twitter. The tool is optimized for short audio clips and may struggle with longer recordings. AI or Not audio detector [374] allows users to upload content or provide URLs, analyzing the input to determine the likelihood of AI generation. ElevenLabs AI Speech Classifier [383] provides a mechanism to verify the origin of synthetic speech, but it can only detect its own generated voices (audio clip created using ElevenLabs technology).
$\clubsuit$ Multi-modal. As multimodal deepfake threats escalate across industries, many organizations turn first to enterprise-grade, subscription-based platforms. Solutions like Sensity AI [385] and Reality Defender[394] build on large, multimodal neural ensembles to scan video, image, audio, and even metadata at scale, trading transparency for robust protection under custom-priced plans. DuckDuckGoose AI (Phocus)[388] and Sentinel[389] extend similar capabilities via paid web interfaces and per-device licensing, while Get Real Protect [393] offers a cross-platform defense suite—each promising high-throughput detection but requiring significant investment and relying on proprietary models that may lag behind the latest adversarial attacks.
At the other end of the spectrum, a wave of freemium and API-first services democratizes access to deepfake detection. Hive AI’s [386] Deepfake Detection API and Resemble Detect [387] let developers plug lightweight classification and voice-benchmarking endpoints into their own workflows, supplemented by free browser extensions or usage quotas. Attestiv [380] and Deepfake Detector [390] flag suspicious video and audio using generative-AI context checks, while BioID [392] zeroes in on facial authenticity via simple web and API calls. These offerings balance ease of integration and cost, inviting smaller teams and researchers to experiment without large upfront commitments—albeit with limited coverage and a few caveats.
Meanwhile, the open-source and academic communities offer transparency and reproducibility in detection research. Tools like Deepware [379] Scanner provide a free web UI and SDK for spotting visual and auditory anomalies, and the University at Buffalo’s DeepFake-o-meter[391] experiments evolve public web-based detectors with no licensing fees.
4.2 Metrics
Evaluation metrics can help quantify the performance and compare the effectiveness of different models. In LAIM-generated media detection tasks, the metrics are mainly from classification scenarios, as we summarize in Table III.
4.2.1 Common Metrics Across Modalities
The most commonly used measures among different modalities include AUC (Area Under the ROC Curve) [43, 14, 102, 76, 72, 114, 63, 62], F1-score [41, 294, 50, 43, 27], FPR (False Positive Rate) [398, 43, 48, 336, 84], Accuracy [27, 60, 52, 37, 43, 96], Precision [52, 166, 32, 83, 14, 308], Recall [83, 14, 32, 166, 43], Average Precision [254, 116, 260, 43, 399].
4.2.2 Modality-Specific Metrics
$\clubsuit$ Text-Specific Metrics. In the text modality, except for above common metrics, [34, 41] used AvgRec (average recall) for binary classification, which is calculated by averaging the recall scores on human-written texts (HumanRec) and LAIM-generated texts (MachineRec).
$\clubsuit$ Image-Specific Metrics. In the context of LAIM-generated image detection, specific evaluation metrics are employed for tasks such as localization and fairness. For forgery localization [92, 254, 255], IoU (Intersection over Union), or its averaged form (mIoU), is the most widely used metric. It quantifies the degree of overlap between the predicted and ground truth masks, providing a spatially sensitive measure of localization accuracy. In addition to IoU, some studies [256, 92, 96] have adopted pixel-level evaluation metrics such as pixel-level AUC and pixel-level F1 score, which assess discriminative performance and the balance between precision and recall at the pixel level. PBCA (Pixel-wise Binary Classification Accuracy), which treats each pixel as an independent sample to measure classification accuracy, is also used in some works [92, 254]. Fairness evaluation in LAIM-generated image detection involves assessing disparities in performance across demographic groups [116]. Demographic Parity $F_{DP}$ measures group-level prediction rate differences. Max Equalized Odds $F_{MEO}$ captures the largest disparity in true or false positive rates between any two groups, while Equal Odds $F_{EO}$ quantifies disparities between subgroup and overall performance. Overall Accuracy Equality $F_{OAE}$ reports the maximum gap in accuracy among all demographic groups. Individual Fairness $F_{IND}$ evaluates whether similar individuals receive similar model outputs.
$\clubsuit$ Video-Specific Metrics. For video detection tasks, Video-level AUC is used to evaluate model performance in works [267, 120], where each video is treated as a single instance by averaging frame-level predictions to compute a single score for AUC calculation.
$\clubsuit$ Audio-Specific Metrics. In audio detection, EER (Equal Error Rate) remains a common choice [132, 240]. It is the location on a ROC curve where the false positive rate and false negative rate are equal. Generally, biometric systems have higher accuracy when the EER is lower. Additionally, the t-DCF (Tandem Detection Cost Function) was created for the ASVspoof challenge [400] to evaluate spoofing countermeasures (CM) in conjunction with Automatic Speaker Verification (ASV). It is used in  [240] to quantify the trade-off between different types of detection errors and the cost associated with them.
$\clubsuit$ Multimodal-Specific Metrics. In the multimodal setting, there are tasks associated with specific evaluation metrics tailored to their objectives. For instance, in [14, 230, 229], the framework involves both image and text grounding, where the detector is required to localize manipulated regions by predicting bounding boxes on the image and identifying corresponding forged segments in the text. For image grounding, the mIoU is used to measure the average overlap between predicted and true bounding boxes. In text grounding, considering the class imbalance scenario that manipulated tokens are much fewer than original tokens, Precision, Recall, and F1 Score are used as evaluation metrics for the manipulated text token grounding.
More recently, researchers have begun using MLLMs (Multimodal Large Language Models), such as GPT-4o and LLaMA2, for LAIM-generated image detection and explanation tasks [227, 228]. These models go beyond binary classification by providing natural language explanations that justify whether an image is real or fake. For evaluating such explanation tasks, metrics like CSS (Cosine Semantic Similarity) and ROUGE (Recall-Oriented Understudy for Gisting Evaluation) are utilized in  [227] and  [228], respectively. CSS measures the semantic alignment between the generated and reference explanations using high-dimensional embeddings, while ROUGE [401] quantifies the overall accuracy of the generated explanations.